help me. please!
a highway patrolman spots a speeding car. he clocks it at 70 mph and takes after it .05 mile behind. if the patrolman travels at an average rate of 90 mph, how long before he over takes he car?

Answer by scott8148(6628)

the rate that the patrolman closes the distance is the difference in the speeds, 90-70 or 20 mph

the distance is .05

r*t=d __ t=d/r __ t=.05/20 __ t=.0025 hr

.0025*60=.15 min __ .15*60=9 sec